OpenStudy (anonymous):

If t varies as v, and t = 2 and 4/7 when v = 13/14 , find v when t = 2 and 1/4.

OpenStudy (anonymous):

((13/14) / (2+(4/7))) * (2+(1/4)) = ((13/14) / ((14/7)+(4/7))) * ((8/4)+(1/4)) = ((13/14) / ((14+4)/7)) * ((8+1)/4) = ((13/14) / (18/7)) * (9/4) = (13/14) * (7/18) * (9/4) = (13*7*9) / (14*18*4) = 819 / 1008 = 13 / 16 = 0.8125

OpenStudy (anonymous):

thank you!!!